OpenStudy (anonymous):

9. Jeremy shuddered _______ when he smacked his elbow on his desk than when he hit his broken foot on the desk. Which word or phrase best completes the sentence above? painfulliest painfullier more painfully most painfully

OpenStudy (anonymous):

c

OpenStudy (anonymous):

C and how can you break a FOOT on a DESK?!!?

OpenStudy (anonymous):

haha bad luck i guess

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

gj

OpenStudy (anonymous):

last one and i pass

OpenStudy (anonymous):

wait Session Score: (7/9) 77.8% Overall Score: (16/29) 55.2%

OpenStudy (anonymous):

C yeah i agree sorry i was helping someone else :p

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

School sucks man, the real question should be \[\eta \le \int\limits_{\sqrt[90]{45}}^{12}\] When will you use this? a) never b) one day at the beach c) on the toilet